# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

### v3

@@ 3.3 Fisher Geometry and Implicit Trust Region
Let $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$ and $\tilde{p}_{\theta}(i;u)=\frac{e^{u_{i}}}{\sum_{j}e^{u_{j}}}$. Let $u^{\star}$ satisfy $\tilde{p}_{\theta}(\cdot;u^{\star})=q(\cdot)$ and define the $q$-centered logits $\delta_{i}=u_{i}-\sum_{j}q_{j}u_{j}$. Then, as $u\to u^{\star}$,
$\mathrm{KL}\!\big(q\,\|\,\tilde{p}_{\theta}(\cdot;u)\big)\;=\;\tfrac{1}{2}\,\delta^{\top}\!\big(\mathrm{Diag}(q)-qq^{\top}\big)\,\delta\;+\;o(\|\delta\|^{2})\;=\;\frac{1}{2\tau^{2}}\,\mathrm{Var}_{q}\!\big[s-s^{\mathrm{ref}}\big]\;+\;o(\cdot),$
where the variance is taken with respect to $q$ and the quadratic form uses the softmax Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ (which is invariant to adding constants to all logits).
@@
ADPO performs optimization in the logit space centered at the reference policy: the coordinate origin is fixed at $s^{\mathrm{ref}}$, so updates are measured as relative log-odds $(s-s^{\mathrm{ref}})$. Near $p\approx q$, the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” around the reference anchor. This contrasts with:
  • Knowledge Distillation (KD): Minimizes $\mathrm{KL}(q\|p_{\theta})$ in the Euclidean probability simplex without a fixed geometric center, making it sensitive to teacher noise.
@@
  • ADPO: Anchoring logits to $s^{\mathrm{ref}}$ yields a distribution-space analogue of TRPO with a fixed geometric center, providing an implicit trust region without explicit constraints.
Figure 1: ADPO’s anchored trust region. The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” in the distribution space centered at the reference anchor. The black trajectory shows the optimization path descending toward the optimal policy, with the anchored coordinate system providing an implicit trust region that prevents large distributional drift. This geometric structure holds for both fixed anchors (stable centers for distillation) and dynamic anchors (adaptive regions for exploration).
Anchor flexibility and update strategies.
Crucially, Lemma 3.3 holds regardless of anchor choice or update strategy. The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ automatically constrains optimization to a local quadratic region, whether the anchor is:
  • Fixed anchor: A reference $\pi_{\text{ref}}$ pre-trained and frozen throughout training, providing a stable geometric center for consistent target tracking;
  • Dynamic anchor: A policy $\pi_{\theta_{\text{old}}}$ that periodically updates to the current policy (every $K$ epochs via $\pi_{\text{ref}}\leftarrow\pi_{\theta}$), providing an adaptive trust region similar to TRPO’s old-policy mechanism;
@@
Our experiments (Section 5) reveal task-dependent trade-offs: dynamic anchoring excels at online exploration in noisy environments (adapting to evolving policies), while fixed anchoring excels at offline distillation from teachers (maintaining stable targets). Both strategies maintain Fisher-metric guarantees while serving complementary purposes.
@@ 5.1 $2\times 2$ Base Comparison Across Scenarios
Table 1: WinMass across 12 scenarios (random baseline = 1/P = 0.25 for P=4). Values are mean over 10 random seeds (std << 0.05 for most entries; 95% CI and Wilcoxon p-values in Appendix). Bold: best method. Underline: best among pairwise. ”Best Listwise” shows the highest-performing variant among ADPO Listwise-Raw/KDE/KDE-Rank, with superscript: $^{\text{R}}$=Raw, $^{\text{K}}$=KDE, $^{\text{KR}}$=KDE-Rank. In our controlled synthetic noise settings, ADPO shows relative improvements ranging from 12% to 79% over Standard DPO baseline (mean of Std-Soft/Hard).
@@
Contrary to conventional wisdom, hard labels dominate under heavy noise (Table 2). Under Heavy Noise-Heavy, ADPO-Hard achieves 0.770 vs. ADPO-Soft’s 0.702 (+9.7%). However, soft labels excel under distribution shift (Light: 0.841 vs. 0.794, +5.9%) and moderate adversarial scenarios. Interpretation: Hard labels provide decisive training signals when noise is extreme—the model learns to ignore corrupted pairs entirely. Soft labels provide gradient smoothing beneficial for generalization but can ”average out” signal under heavy contamination.
@@ 5.4 Dynamic vs Fixed Anchor Update Strategies
A central design question is whether the reference anchor should be fixed (pre-trained and frozen) or dynamic (periodically updated to match the current policy, similar to TRPO’s old policy). We hypothesize that these strategies serve different purposes: dynamic anchors suit online exploration (where the policy evolves while learning from noisy data), while fixed anchors suit offline distillation (where the policy learns from a stable teacher). Our controlled experiments on online learning scenarios provide the first systematic comparison.
@@ 5.6 Hyperparameter Sensitivity
To assess the robustness of our method to hyperparameter choices, we perform a sensitivity analysis on the temperature parameters $\beta_{r}$ and $\tau$. Figure 5 shows the results. Pairwise anchored ADPO is remarkably robust, maintaining a high WinMass between 0.61–0.62 across all tested combinations of $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$. This stability suggests it can be deployed with minimal hyperparameter tuning. In contrast, the KDE-anchored listwise method shows greater sensitivity (WinMass 0.37–0.55), indicating that it benefits from careful tuning.
@@ 6 Discussion
In our synthetic noise experiments, hard labels dominate under heavy noise (8/12 scenarios) because they provide binary gradients that either fully update or ignore corrupted pairs—effectively an implicit outlier detector when combined with anchoring. Soft labels with $q_{ij}\approx 0.5$ produce weak gradients that ”average out” signal. However, under distribution shift, soft labels preserve calibration gradients, allowing the model to adapt confidence. Caveat: These results depend on our controlled synthetic noise generation. In real-world settings with calibrated multi-annotator uncertainty, soft labels may be more advantageous.
@@ 7 Limitations and Future Work
  • Reference update strategies: We use frozen references. Comparing frozen vs. EMA slow-update ($\tau_{\text{ema}}=0.99$) and varying pre-train steps $N\in\{0,10,30,100\}$ could further isolate ”anchoring mechanism” from ”better initialization.”

### v4

@@ 3.4 Fisher Geometry and Implicit Trust Region
Let $u=(s-s^{\mathrm{ref}})/\tau$ and $\tilde{p}(u)=\mathrm{softmax}(u)$. Let $u^{\star}=\log q+c\mathbf{1}$ so that $\tilde{p}(u^{\star})=q$ (the softmax is shift-invariant). Define $\delta=u-u^{\star}$ and center it by $q$ as $\bar{\delta}=\delta-\langle q,\delta\rangle\mathbf{1}$. Then, as $u\to u^{\star}$,
$\mathrm{KL}\!\big(q\,\|\,\tilde{p}(u)\big)\;=\;\tfrac{1}{2}\,\bar{\delta}^{\top}\!\big(\mathrm{Diag}(q)-qq^{\top}\big)\,\bar{\delta}\;+\;o(\|\bar{\delta}\|^{2})\;=\;\frac{1}{2\tau^{2}}\,\mathrm{Var}_{i\sim q}\!\big[(s_{i}-s_{i}^{\mathrm{ref}})-\tau\log q_{i}\big]\;+\;o(\cdot),$
where the variance is taken with respect to $q$, measuring the deviation of the anchored log-odds from the target $q$. The quadratic form uses the softmax Fisher metric $\mathrm{Diag}(q)-qq^{\top}$, which has a null-space along $\mathbf{1}$ (reflecting shift-invariance).
@@
Corollary 3.4 (Local bounds).
Let $\mu(q)>0$ denote the smallest nonzero eigenvalue of $\mathrm{Diag}(q)-qq^{\top}$ (the null-space is along $\mathbf{1}$). Then locally around $u^{\star}$,
$\mathrm{KL}(q\|\tilde{p}(u))\;\geq\;\frac{\mu(q)}{2}\,\|\bar{\delta}\|^{2},$
where $\bar{\delta}=\delta-\langle q,\delta\rangle\mathbf{1}$ is the $q$-centered deviation. By Pinsker’s inequality, $\|\tilde{p}(u)-q\|_{1}\leq\sqrt{2\mathrm{KL}(q\|\tilde{p}(u))}$, providing a distributional deviation bound.
On the smallest nonzero eigenvalue: For $K=2$, $\mu(q)=2q(1-q)$. For $K>2$, $\mu(q)$ depends on the full spectrum of $q$; a uniform, distribution-independent lower bound is generally loose. When $\min_{i}q_{i}\geq\varepsilon>0$, the smallest nonzero eigenvalue satisfies $\mu(q)\geq\varepsilon$. Proof sketch: For any $v\in\mathbf{1}^{\perp}$, the Rayleigh quotient gives $v^{\top}(\mathrm{Diag}(q)-qq^{\top})v=\sum_{i}q_{i}v_{i}^{2}-(\sum_{i}q_{i}v_{i})^{2}=\mathrm{Var}_{q}[v]\geq\varepsilon\|v\|^{2}$ since $\sum_{i}v_{i}=0$ and the variance of $v$ under any distribution bounded below by $\varepsilon$ is at least $\varepsilon\|v\|^{2}$.
@@
ADPO performs optimization in the anchored $u$-coordinate system: the reference $\pi_{\mathrm{ref}}$ defines the coordinate origin via $u=(s-s^{\mathrm{ref}})/\tau$, but the center of the local quadratic bowl is at $u^{\star}$ satisfying $\tilde{p}_{\theta}(\cdot;u^{\star})=q(\cdot)$, not at the anchor itself. The local curvature at $u^{\star}$ is governed by the softmax Fisher metric $\mathrm{Diag}(q)-qq^{\top}$, which depends only on the target $q$. Thus: the anchor determines the coordinate system, while the target $q$ determines both the minimizer $u^{\star}$ and the local geometry. This contrasts with:
  • Knowledge Distillation (KD): Minimizes $\mathrm{KL}(q\|p_{\theta})$ in the Euclidean probability simplex. While the Fisher curvature exists at any point, there is no fixed reference anchor, so the geometry lacks a stable coordinate origin.
@@
  • ADPO: Anchoring logits to $s^{\mathrm{ref}}$ yields a distribution-space analogue of TRPO with a fixed coordinate origin, inducing an implicit trust region without explicit constraints.
Figure 1: ADPO’s anchored trust region. The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” in the distribution space, expressed in anchored coordinates $u=(s-s^{\mathrm{ref}})/\tau$; the quadratic well is centered at $u^{\star}$ (i.e., at the target $q$), not at the anchor itself. The green trajectory shows the optimization path descending toward the optimal policy, with the anchored coordinate system providing an implicit trust region that prevents large distributional drift. This geometric structure holds for both fixed anchors (stable centers for distillation) and dynamic anchors (adaptive regions for exploration).
3.5 Geometric Interpretation: KL-Mirror Descent in $u$-Space
Recall $A(u)=\log\!\sum_{j}\exp(u_{j})$ and $\tilde{p}(u)=\nabla A(u)=\mathrm{softmax}(u)$, and our anchored objective $\mathcal{L}(u)=A(u)-\langle q,u\rangle=\mathrm{KL}\!\big(q\;\|\;\tilde{p}(u)\big)+H(q)$. Let $u_{t}$ be the current anchored logits and $g_{t}=\nabla\mathcal{L}(u_{t})=\tilde{p}(u_{t})-q$.
Theorem 3.5 (KL-mirror descent / Bregman-prox step in $u$-space).
If one were to directly optimize $\mathcal{L}$ over the anchored logits $u$ (rather than shared parameters $\theta$), then for any step size $\eta>0$, the one-step update can be written equivalently as the solution of the following proximal subproblem:
$u_{t+1}\ =\ \arg\min_{u}\ \Big\langle g_{t},\ u\Big\rangle\ +\ \frac{1}{\eta}\,\mathrm{KL}\!\Big(\tilde{p}(u_{t})\,\big\|\,\tilde{p}(u)\Big).$ (11)
Equivalently, since $D_{A}(u\|u_{t})=A(u)-A(u_{t})-\langle\nabla A(u_{t}),u-u_{t}\rangle=\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u))$, (11) is a standard mirror descent / Bregman-prox step with generator $A$.
Proof.
Linearize $\mathcal{L}$ at $u_{t}$: $\mathcal{L}(u)\approx\mathcal{L}(u_{t})+\langle g_{t},u-u_{t}\rangle$. A mirror descent step with generator $A$ minimizes $\langle g_{t},u\rangle+\tfrac{1}{\eta}D_{A}(u\|u_{t})$. Using $D_{A}(u\|u_{t})=\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u))$ for the softmax family yields (11). ∎∎
Dual form in $p$-space. Via the Legendre map $p=\nabla A(u)$, the step (11) is equivalent to the entropic proximal in probability space
$\tilde{p}_{t+1}=\arg\min_{p\in\Delta}\ \langle g_{t},p\rangle+\tfrac{1}{\eta}\mathrm{KL}\!\big(p\,\|\,\tilde{p}_{t}\big),$
whose solution is the multiplicative-weights update $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta g_{t})$.
Multiplicative-weights view.
The solution of (11) is equivalent to the multiplicative-weights update in the anchored distribution:
$\tilde{p}_{t+1}\ \propto\ \tilde{p}_{t}\ \odot\ \exp\!\big(-\eta\,g_{t}\big)\ =\ \tilde{p}_{t}\ \odot\ \exp\!\big(-\eta\,(\tilde{p}_{t}-q)\big),$ (12)
followed by renormalization. This exhibits ADPO as a KL-trust-region step in distribution space.
Step-size $\eta$ vs temperature $\tau$ (heuristic).
Anchored coordinates $u=(s-s^{\mathrm{ref}})/\tau$ rescale curvature by $1/\tau^{2}$ (Lemma 3.3). To maintain comparable “trust-region tightness” in the unanchored score space $s$, one may heuristically choose
[…]
This is a practical guideline; a rigorous derivation would require accounting for the chain rule $\partial u/\partial\theta$ when optimizing over shared parameters $\theta$.
3.6 Dynamic Anchor Stability
Crucially, Lemma 3.3 holds regardless of anchor choice. The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ automatically constrains optimization to a local quadratic region. While the coordinate $u=(s-s^{\mathrm{ref}})/\tau$ depends on the anchor, the local quadratic form at the optimum only depends on $q$; thus the curvature (Fisher metric) is anchor-agnostic, albeit the optimization objective itself changes with the anchor. Common anchor choices include:
  • Fixed anchor: A reference $\pi_{\text{ref}}$ pre-trained and frozen throughout training, providing a stable coordinate origin for consistent target tracking;
  • Dynamic anchor: A policy $\pi_{\theta_{\text{old}}}$ that periodically updates to the current policy (every $K$ epochs via $\pi_{\text{ref}}\leftarrow\pi_{\theta}$), providing an adaptive coordinate system similar to TRPO’s old-policy mechanism;
@@ 3.6 Dynamic Anchor Stability
For dynamic anchors, we characterize the monotonic improvement within each fixed-anchor phase:
Theorem 3.6 (Fixed-anchor-phase monotonic descent in $u$-space).
Consider a training phase where the anchor $\pi_{\mathrm{ref}}$ is held fixed. Define $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ where $A(u)=\log\sum_{j}e^{u_{j}}$ and $u=(s-s^{\mathrm{ref}})/\tau$ are the anchored logits. When directly optimizing $\mathcal{L}$ over $u$ (e.g., tabular setting or each group having independent degrees of freedom), $\mathcal{L}$ is strictly convex on the subspace $\mathbf{1}^{\perp}$ (orthogonal to the all-ones vector) with Hessian $\nabla^{2}\mathcal{L}(u)=\mathrm{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$, which is positive semidefinite on $\mathbf{1}^{\perp}$ with one zero eigenvalue along $\mathbf{1}$.
Strong convexity (local or bounded-domain):
  • Local: In a neighborhood of $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$), the Hessian $\nabla^{2}\mathcal{L}(u)=\mathrm{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$ approaches $\mathrm{Diag}(q)-qq^{\top}$ up to $o(1)$, so there exists $\mu>0$ (depending on $q$) such that $\mathcal{L}$ is $\mu$-strongly convex on $\mathbf{1}^{\perp}$ in this neighborhood.
  • Bounded domain: If the optimization is constrained to a domain where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$ for all $i$ and some $\varepsilon>0$ (i.e., the current anchored distribution $\tilde{p}_{\theta}(u)$ is bounded away from the probability simplex boundary), then $\nabla^{2}\mathcal{L}(u)\succeq\varepsilon\,P_{\mathbf{1}^{\perp}}$ globally in this domain (one may take $c(\varepsilon)=\varepsilon$).
Under either condition, standard gradient descent with appropriate step size $\eta\in(0,1/L]$ (where $L$ is a local smoothness constant) yields linear (geometric) convergence:
$\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\;\leq\;(1-\mu\eta)\,(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star})),$
implying monotonic decrease in $\mathrm{KL}(q\|\tilde{p}_{\theta})$ within each fixed-anchor phase.
Note: When the anchor is updated (e.g., $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$), the objective $\tilde{p}_{\theta}$ is re-normalized with the new anchor, so the loss function changes. We do not claim monotonicity across anchor updates; instead, each fixed-anchor phase provides stable descent, and the anchor update adapts the coordinate system for the next phase.
Caveat for parametrized models: In practice, the anchored logits $u$ are derived from shared parameters $\theta$ (e.g., neural networks), and the optimization problem over $\theta$ is typically non-convex. The above results provide geometric intuition for the loss landscape in the anchored distribution space but do not constitute convergence guarantees for parameterized training.
Proof sketch.
The log-partition function $A(u)$ is strictly convex (its Hessian is the softmax Fisher metric $\mathrm{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$, which is positive semidefinite with null-space along $\mathbf{1}$). The Hessian depends on the current $\tilde{p}_{\theta}(u)$, not directly on $q$. However, near $u^{\star}$ where $\tilde{p}_{\theta}(u^{\star})=q$, it coincides with $\mathrm{Diag}(q)-qq^{\top}$ up to higher-order terms, yielding local strong convexity. Alternatively, if $\tilde{p}_{\theta}(u)$ is constrained to $\varepsilon,1-\varepsilon]^{K}$, the minimum eigenvalue is uniformly bounded below. Standard strongly convex optimization theory then guarantees linear convergence. When the anchor changes, $u$ is redefined, so objectives from different phases are not directly comparable. ∎∎
Implication. This theorem explains why dynamic anchors (Section 5.4) maintain stability during online exploration: within each fixed-anchor phase, the implicit Fisher-metric trust region ensures monotonic descent; across anchor updates, the coordinate system adapts to track policy evolution, similar to TRPO’s mechanism. Our experiments (Section 5) reveal task-dependent trade-offs: dynamic anchoring excels at online exploration in noisy environments (adapting coordinates to evolving policies), while fixed anchoring excels at offline distillation from teachers (maintaining a stable coordinate origin throughout). Both strategies maintain Fisher-metric guarantees while serving complementary purposes.
@@ 4 Experimental Setup
Evaluation metrics.
We use three primary metrics to assess preference learning quality:
  • WinMass: For each context $x$ with candidate set $S_{x}=\{1,\ldots,P\}$ and ground-truth ranking (higher reward is better), WinMass measures the total probability mass assigned to the top item under the anchored distribution: $\mathrm{WinMass}=\mathbb{E}_{x}[\tilde{p}_{\theta}(\operatorname*{arg\,max}_{i}R_{i}^{\star}\mid S_{x})]$. Higher is better; random baseline = $1/P$.
  • NDCG@k: Normalized Discounted Cumulative Gain at position $k$, a standard ranking metric that rewards placing high-reward items at top positions with logarithmic discounting. Higher is better; range $[0,1]$.
  • KL divergence: For distillation tasks, we measure $\mathrm{KL}(\pi_{\mathrm{teacher}}\|\pi_{\mathrm{student}})$, the expected KL divergence from teacher to student policy over states and actions. Lower is better, indicating closer alignment with the teacher.
@@ 5.1 $2\times 2$ Base Comparison Across Scenarios
Table 3: WinMass across 12 scenarios (random baseline = 1/P = 0.25 for P=4). Values are mean over 10 random seeds (std << 0.05 for most entries; 95% CI and Wilcoxon p-values in Appendix). Bold: best method. Underline: best among pairwise. "Best Listwise" shows the highest-performing variant among ADPO Listwise-Raw/KDE/KDE-Rank, with superscript: $^{\text{R}}$=Raw, $^{\text{K}}$=KDE, $^{\text{KR}}$=KDE-Rank. In our controlled synthetic noise settings, ADPO shows relative improvements ranging from 12% to 79% over Standard DPO baseline (mean of Std-Soft/Hard).
@@
Contrary to conventional wisdom, hard labels dominate under heavy noise (Table 4). Under Heavy Noise-Heavy, ADPO-Hard achieves 0.770 vs. ADPO-Soft’s 0.702 (+9.7%). However, soft labels excel under distribution shift (Light: 0.841 vs. 0.794, +5.9%) and moderate adversarial scenarios. Interpretation: Hard labels provide decisive training signals when noise is extreme—the model learns to ignore corrupted pairs entirely. Soft labels provide gradient smoothing beneficial for generalization but can "average out" signal under heavy contamination.
@@ 5.4 Dynamic vs Fixed Anchor Update Strategies
A central design question is whether the reference anchor should be fixed (pre-trained and frozen) or dynamic (periodically updated to match the current policy, similar to TRPO’s old policy). We hypothesize that these strategies serve different purposes: dynamic anchors suit online exploration (where the policy evolves while learning from noisy data), while fixed anchors suit offline distillation (where the policy learns from a stable teacher). Our controlled experiments on online learning scenarios provide a systematic empirical comparison.
@@ 5.6 Hyperparameter Sensitivity
To assess the robustness of our method to hyperparameter choices, we perform a sensitivity analysis on the temperature parameters $\beta_{r}$ and $\tau$. Figure˜5 shows the results. Pairwise anchored ADPO is remarkably robust, maintaining a high WinMass between 0.61–0.62 across all tested combinations of $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$. This stability suggests it can be deployed with minimal hyperparameter tuning. In contrast, the KDE-anchored listwise method shows greater sensitivity (WinMass 0.37–0.55), indicating that it benefits from careful tuning.
@@ 6 Discussion
In our synthetic noise experiments, hard labels dominate under heavy noise (8/12 scenarios) because they provide binary gradients that either fully update or ignore corrupted pairs—effectively an implicit outlier detector when combined with anchoring. Soft labels with $q_{ij}\approx 0.5$ produce weak gradients that "average out" signal. However, under distribution shift, soft labels preserve calibration gradients, allowing the model to adapt confidence. Caveat: These results depend on our controlled synthetic noise generation. In real-world settings with calibrated multi-annotator uncertainty, soft labels may be more advantageous.
@@ 7 Limitations and Future Work
  • Reference update strategies: We use frozen references. Comparing frozen vs. EMA slow-update ($\tau_{\text{ema}}=0.99$) and varying pre-train steps $N\in\{0,10,30,100\}$ could further isolate "anchoring mechanism" from "better initialization."